aug15: lbw polB drifting
Proj A2812 data
19aug15: 900 secs of total power data while
31aug15: 2 hours of total power data while
sitting. warmup then cooldown dewar.
17,18sep15: total power data after dewar worked
22sep15: total pwr vs time as we warm the dewar a
Project a2812 reported unstable
baselines in their lbw data taken in august. I looked at their data.
We then took some total power data on 19,31aug15, 17,18,22 sep15.
Their setup is:
I plotted up the total power vs
time for the data they took in aug15 (.ps) (.pdf)
- lbw linear pol
- redshifted H1.. 1360-1420
- interim correlator
- sbc1 polA H1
- sbc2 polB H1
- sbc3 PolA OH
- sbc4 polB OH
- 1 second dumps
- each frame contains the total power vs time for a nights run.
- black is pola, red is polB
- They usually do on, off position switching with a cal
- 300 secs on source
- 300 secs off source
- then fire the cal
- Each narrow spike is the cal firing
- between the cal spikes, you should see a power variation (on
za change) and then a similar repetition (on off source over
same za range).
- The vertical scale is linear in power
- You can see the red trace (polB) drifting during most night's
19aug15:Total power data taken while sitting:
On 19aug15 the stability of lbw was monitored using the interim
correlator. The setup was:
The plots show the total power vs time
while sitting at az=285,za=11 (.ps) (.pdf)
- lbw,linear pol, cfr=1405 MHz
- 1370 hipass filter and 1360 to 1470 band pass filter
- (i also tried it with the default 1100 to 1800 MHz filter
with similar results).
- interim correlator,
- 25 MHz polA,B in each of 4 sbc (to make sure the problem
wasn't in the iflo or interim correlator.
- 1 second dumps for 910 seconds. telescope sitting at az=285,
- Black is polA, red is PolB
- There are 4 polA traces and 4 polB traces.
- the traces within each set of 4 are identical
- This means that the problem is before the 4 mixers
downstairs that splits up the band.
- the bumps at 310 and 400 seconds are sources drifting through
- Each trace was normalized to it's median value and then 1 was
subtracted.. so the vertical scale is in units of Tsys.
- You can see that polB is drifting around by up to 1% of Tsys.
150831: totPwr data taken while dewar
warmed then cooled.
On 31aug15 we looked at the stability of lbw
while we let the dewar warm up a little.
the setup was:
The plots show the
total power vs time (.ps) (.pdf):
- the telescope was at: az=284.7, za=11
- lbw was setup with cfr=1405 Mhz, 1370hipass and 1360-1470 Mhz
- interim correlator took 25Mhz centered at 1405 Mhz using 4
separate boards, and 1 second dumps
- the extra boards were to check if there was an instability
in the if/lo or interim cor.
- all 4 were identical..
- The total power for each 1second dump was computed over the
25 Mhz bw.
- i averaged the 4 separate boards, since they showed no
- Black is polA, red is PolB. The vertical scale is Tsys (i
removed the average)
- The green line is the 16K stage dewar temp (divided by 100).
- The green line shows when we started warming up the dewar,
cooling it back down, and the warming it up again.
- PolA Tsys is stable during these changes
- PolB reacts to changes in the 16K stage.
- around 9.3 hours, the 16K stage went from 20K to 28Kthen
back down to 20k.
- polB power output power went from tsys to 1.38 Tsys.
- It looks like polB is reacting the changes in the 16K stage.
150917,18 : Data taken after dewar worked
The dewar was taken offline and worked on during
sept. The voltages for the compressors and refrigerators were also
16sep15 (this stopped the problem of ratcheting in the dewar.
Data was taken on 17 and 18sep15 to again check
the stability of lbw.
The setup was:
- lbw center freq of 1405, 1370 MHz hipass filter installed as
well as the 1360-1470 MHz filter
- interim correlator using 25MHz bandpass, 1 second dumps, 4
- Total power was computed for each 1 second dump.
The plots show the total power vs time for
the 2 days (.ps) (.pdf)
- black in polA, red is polb
- Top: 17sep15
- the narrow spikes are from punta salinas transmitting in all
- PolB starts to drift up at 16.2 hours.
- This may be polarized power (north polar spur) or it could
be a problems with lbw polB
- Bottom: 18sep15
- both traces follow one another
- The rise at 11.8 hours is the sun sidelobes coming into the
- For the measurements on both of these days, the dewar
temperature was stable.
150922:totpwr data while we warm up dewar
The data from 18sep15 showed that lbw polA and
polB tracked one another for over an hour. During this observation
the dewar temperature was stable. Back on 31aug15 we saw that as
we warmed up the dewar a little, polB signal increased rapidly.
On 22sep15 we turned off the lbw refrigerator
for a few minutes while we took some total power data.
The setup was:
The plots show the total power vs
time for the 500 sec observation (.ps) (.pdf)
- the telescope was sitting at az=21.5, za=1.8998
- lbw linear pol, center freq 1405Mhz, 1370 hipass and
1360-1470 filters in.
- interim correlator took 25 Mhz bw of data using 4 boards,
9 level sampling, 1 second integrations.
- we computed the total power from the 25Mhz band.
- We turned off the lbw refrigerator at 9:57:49 (9.96361 hour).
- Black is pola, red is polB. The vertical scale is in units of
Tsys ( 1 has been removed).
- green is the the 16 K dewar temperature. The scale is
degK/100. So the temp started at 15.5K.this data is measured
once every 32 seconds.
- Top plot: 499 seconds of data.
- You can see polB power head as soon as the Tsys starts to
- We have undersampled the 16k Temp. The refrigerator was
turned off at the blue dashed line.
- Bottom: blow showing totpwr vs seconds from refrig turn off.
- The polB started to rise about 3 seconds after the shutdown
(each + for the polA,B is a 1 second sample).
- The fall in the 16k temp was when the refrigerator was
turned back on.
- a2812 reported unstable baselines for their aug15 data.
- The problem was with polB (not polA as they mentioned..)
- data taken on 19aug15 showed that polB is drifting by up to 1%
- The problem occurred with the 1100-1800 MHz filter as well as
the 1370 hipass and1360-1470 band pass filter
- so the problem is not in the filter box.
- On 31aug15 we warmed up the dewar a little and the power level
for polB shot way aup
- After working on the dewar in sept, data was taken on
- On 17sep15 there was a slow rise in polB (relative to polA)
- On 18sep15 both pols tracked on another.
- The 22sep15 data shows that polB is still extremely
sensitive to 16K dewar temperature variations
- The signal increased 35% in 50 seconds.. as the 16k stage
went from 15K to 23K.
- We still have a problem with polB.